Back to the NIKA2 Pool page

                                                             _.-=;~ /_
                                                          _-~   '     ;.
                                                      _.-~     '   .-~-~`-._
                                                _.--~~:.             --.____88
                              ____.........--~~~. .' .  .        _..-------~~
                     _..--~~~~               .' .'             ,'
                 _.-~                        .       .     ` ,'
               .'                                    :.    ./
             .:     ,/          `                   ::.   ,'
           .:'     ,(            ;.                ::. ,-'
          .'     ./'.`.     . . /:::._______.... _/:.o/
         /     ./'. . .)  . _.,'               `88;?88|
       ,'  . .,/'._,-~ /_.o8P'                  88P ?8b
    _,'' . .,/',-~    d888P'                    88'  88|
 _.'~  . .,:oP'        ?88b              _..--- 88.--'8b.--..__
:     ...' 88o __,------.88o ...__..._.=~- .    `~~   `~~      ~-._ NIKA2 _.
`.;;;:='    ~~            ~~~                ~-    -       -   -

Daily Reports

November 20, (Tuesday)

Laurence Perotto, Jean-François Lestrade, Stefano Berta, Charlène Lefèvre, Marina Ricci and Bilal Ladjelate arrived at the telescope. Bad weather.

November 21, (Wednesday)

NIKA2 Duopomp stopped around 6AM. JLS noticed it and restarted it manually. Short warm up but no apparent harm. Bad weather, still.

November 22, (Thursday)

Bad weather, again. We took 10 minutes to play in the snow and have a snowball fight.

November 23, (Friday)

This morning, Alain is testing a new new version of the acquisition (V3?). We are still not observing.
Damned this is again a weather unlucky run! The difference between acquisition V1 and V2 is a complete re-writing of the code, so Alain works currently on V2.something (only "small" modifications)... ;-) You may have a good observing window tonight, so good luck ! SL. -> I'm out of lucky charm! BL.

The operators, JLS and SA went on the plateform to clean the telescope from ice. But there is still ice on the subreflector. We wait until it melts and we'll start observations.

17:22 WE START OBSERVATIONS! For the run, we will use by default the acquisition V1, and PIIC/RTA alongside, comparing their results. A chunk of ice might still be around, but we can start observe as long as we don't go over 60°. We start with calibrations, many foci, a beammap, and then we'll see if we can move on science observations.

November 24, (Saturday)

Night Shift: Laurence and Marina After some cailbration scans and a beammap, we observe projects 094-18 and the LPSZ until 3AM, and continue with LPDF in good weather conditions. ~3AM, an cryostat alarm rang, signaling the loss of control of the pods levels of NIKA2. The AoD decides to continue observations despite this alarm, and Martino confirmed in the morning that as long as the drift of the pods stays within 1mm, there is no worry to have.

Morning shift: Charlène and Jean-François We continue observations of the LPDF, in excellent and extremely stable weather conditions. After that we continue with project 005-18.

November 25, (Sunday)

During whole sunday, the atmosphere is too bad to observe. 20181126_024151.jpg

Bilal took the opportunity to bring the observers to the cabin and show them NIKA2 at 16h30 (cf. bump in NIKA2 legs levels). 20181125_165420.jpg

November 26, (Monday)

The weather starts improving in the morning, but too much wind, we can't start the antenna before 2PM. Though, during the afternoon shift, while the opacity was excellent (cf figure), 20181126_192640.jpg

November 27, (Tuesday)

The weather kept being stable until the end of the shift, with an exceptionnal opacity and a very good atmosphere stability. We finished the run with LPDF (192-16) and 005-18 observations, all in very good conditions.

DailyReportsNika2Pool201118 (last edited 2018-11-27 09:38:55 by ladjelate)